OpenStudy (anonymous):

Y=2e^x Can someone explain how they get the answer so I'm not brain dead in class?

jimthompson5910 (jim_thompson5910):

What are the full instructions?

OpenStudy (anonymous):

To find the derivative

OpenStudy (fibonaccichick666):

so, what is the derivative of e^x?

OpenStudy (anonymous):

I have no clue I missed that day of class because I was sick and he told me to get notes from a friend but the homework's due tomorrow

OpenStudy (fibonaccichick666):

ok, well, lucky for you, the derivative of e^x is the easiest in existence :)

OpenStudy (fibonaccichick666):

It is just, e^x. It stays the same

OpenStudy (anonymous):

Really?

OpenStudy (fibonaccichick666):

yup

OpenStudy (fibonaccichick666):

it's special

OpenStudy (anonymous):

So does that mean by chance the derivative of this equation the same equation?

OpenStudy (fibonaccichick666):

yep

jimthompson5910 (jim_thompson5910):

yep that's how the number "e" is set up. It is its own derivative if y = e^x, then dy/dx = e^x

OpenStudy (anonymous):

Thank you so much

OpenStudy (fibonaccichick666):

we could prove it... but it is way easier to just take that as given

OpenStudy (fibonaccichick666):

np, good choice of day to miss class

OpenStudy (anonymous):

So if any equation has e it's just the same equation as the derivative or are their special rules?

OpenStudy (fibonaccichick666):

ah, so the chain rule comes into play if you have a function of x instead of just x as an exponent

OpenStudy (anonymous):

Oh okay thank you both so much

OpenStudy (fibonaccichick666):

so, let's say you have \[e^{4x^2+2}\] the derivative of f(x) which I will call u, is \[u'=8x\] so then you need to derive \[e^u\]which equals\[e^u du=e^{4x^2+2}(8x)\]